(6%) Problem 10: Two point charges Q1-Q2-+1.8 nC are fixed symmetrically on the x-axis at x = ±0.062 m. A point particle of charge Q3 = +4.4 μ.C and mass m 13 mg can move freely along the y-axis If the particle on the y-axis is released from rest at y-0.016 m, what will be its speed, in meters per second, when it reaches y2 0079 m? Consider electric forces only.
Answer #1

10.

8 メー-6.062 USin +6.062)2

using conservation of energy

electric potential energy of Q3 at A = electric potential energy of Q3 at B + kinetic energy

kQ1Q3/r + k Q2 Q3/r = kQ1Q3/r' + k Q2 Q3/r' + (0.5) m v2

k (Q1 + Q2) Q3/r = k (Q1 + Q2) Q3/r' +   (0.5) m v2

(9 x 109) ((1.8 x 10-6)+ (1.8 x 10-6)) (4.4 x 10-6)/0.064 = (9 x 109) ((1.8 x 10-6)+ (1.8 x 10-6)) (4.4 x 10-6)/0.1 +   (0.5) (13 x 10-6) v2

v = 351.24 m/s